OpenStudy (anonymous):

The population of Guatemala in 2000 was 12.7 million. Assuming exponential growth, what would be the size of the population after time t (measured in years after 2000) if the population was 30 million in 2075? Answer (in millions): P(t)=

OpenStudy (anonymous):

can anybody help me

OpenStudy (anonymous):

when you need the size of population

OpenStudy (anonymous):

i already know the size of the population. i need the function

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@robtobey @radar @phi @precal

OpenStudy (phi):

Assuming exponential growth, means assume the equation \[ p(t) = A e^{B t} \] where t is years since 2000, and the population p(t) is in millions of people now you need to figure out what A and B are. use The population of Guatemala in 2000 was 12.7 million. In the year 2000, we have t=0 (zero years from 2000) use those numbers in the equation 12.7 = A e^{B\cdot 0}= A \cdot 1 = A that means A is 12.7 now you have \[ p(t) = 12.7 e^{B t} \] use population was 30 million in 2075 (which means p(t) =30 and t= 75) to find B

OpenStudy (anonymous):

b=.01?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

substitute it dude..

OpenStudy (anonymous):

i got b=.01 i just dont know if its right.

OpenStudy (anonymous):

lols..how come u cant know --' this is basic math u know.. c:

OpenStudy (anonymous):

i think i am mentally challenged

OpenStudy (anonymous):

haha..lols..

OpenStudy (anonymous):

so can u help me

OpenStudy (anonymous):

where you come from dude ?

OpenStudy (anonymous):

thats irrelevant. what is relevant is if u can help me with my problem or not

OpenStudy (anonymous):

hahaha..just answer it..dont be afraid to get wrong answer..omg ! i got wrong answered ! T..T my mum will punish me..please help me ==' lolz..

OpenStudy (anonymous):

@Krishnadas

OpenStudy (anonymous):

i am not scared to get it wrong. it is online hw that i have to get right otherwise it wont get turned in on time

OpenStudy (anonymous):

@Frostbite help him dude :) i have another work to settle ;) @muzzammil.raza : what did u said ?? =='

OpenStudy (frostbite):

Assume we don't do natural exponential form, but on the form \[\LARGE f(x)=b*a ^{x}\] the constants b and a are given by: \[\LARGE a=\sqrt[x _{2}-x _{1}]{\frac{ y _{2} }{ y _{1} }}\] b can be found by using 1 equation with 1 variable (substitute a data set into the equation)

OpenStudy (frostbite):

in other words: \[\LARGE b=\frac{ y _{1} }{ a ^{x _{1}} }=\frac{ y _{2} }{ a ^{x _{1}} }\]

OpenStudy (anonymous):

i got it i am good. thanks for all the help

OpenStudy (anonymous):

wtf ! haha..lolz